# Efficient Frame Extraction: A Novel Approach Through Frame Similarity and Surgical Tool Tracking for Video Segmentation

### v1

@@ III-D2 Two-stream Convolutional Networks
Back to the Figure 1, in the Phase Segmentation module, we trained two distinct X3D models separately. In the last layer of both models, instead of outputting classifications, we modified the layer to produce output probabilities. We then ensembled these probabilities to make classifications.
@@ IV-A Benchmark Dataset
The dataset used in the work is obtained from Robotic Pancreaticoduodenectomy (also known as the Whipple procedure)  [53, 54, 55], which is a key procedure for treating pancreatic cancer. It involves two main phases: Resection and Reconstruction. The Reconstruction phase includes three anastomoses: Hepaticojejunostomy, Pancreaticojejunostomy, and Gastrojejunostomy. In the following, we will discuss the GJ and PJ datasets that are used in this study in more detail.
@@ IV-A1 GJ Dataset
A retrospective case review spanning from 2017 to 2021 was conducted at two prominent referral centers, UT Southwestern Medical Center and University of Pittsburgh Medical Center. During this review, each robotic GJ video was segmented into six specific tasks: 1.1 Stay suture, 1.2 Inner running suture, 1.3 Enterotomy, 2.2 Inner running suture, 3.1 Inner Layer of Connell, and 4.1 Outer Layer of Connell, along with idle time. These videos were annotated by experts using a custom annotation framework (See Figure 7).
@@
Of the 42 videos included, 30 were used for training and 12 for validation, with all frames extracted (6 frames/second) and annotated, resulting in a dataset of 46 9034690346\,90346 903 images for training and 22 6472264722\,64722 647 images for validation.
IV-A2 PJ Dataset
A retrospective review was conducted on recording videos collected from 2011 to 2022 across two quaternary referral centers: UT Southwestern Medical Center and the University of Pittsburgh Medical Center [3]. Out of the 100 videos reviewed, 60 were utilized for model training, 10 for hyperparameter optimization, and 30 for performance testing. Frames were extracted at 6 frames/second and annotated. Each video was divided into tasks, with start and end times recorded for each. Six tasks were annotated per video: 1.1 Anterior Mattress Sutures, 1.2 Tying Mattress Sutures, 2.1 Enterotomy, 3.1 Posterior Duct Sutures, 3.2 Anterior Duct Sutures, and 4.1 Anterior Buttress Sutures.
IV-B Distribution of Tools and Naming Conventions
Figure 8 displays the distribution of the number of frames for each class ID corresponding to the sixteen objects listed in Table II for GJ dataset. We used YOLOV8, as described in Section III-C, to extract and track surgical instruments. Each instrument is segmented into three parts, each assigned a unique ID: the shaft, wrist, and jaw, except for the “laptool”. It is important to note that in our dataset, instrument number two (“instrument2”) is identical to instrument number three (“instrument3”) because they are both needle drivers; therefore, they are combined under the same object category. As a result, the class IDs 3, 4, and 5 are merged into IDs 2, 1, and 0.
Figure 8: Distribution of Class IDs. Total of sixteen objects are tracked.
Since the tools have multiple parts and few tools are identical to others, to avoid confusion in our analysis of tools for object detection, we define the following naming conventions:
  • “One Object”: Refers to tracking part 1 (the jaws) of the Needle Driver located on the right-hand side of the screen.
  • “Two Objects”: Refers to tracking part 1 (the jaws) of each of the two Needle Drivers-one located on the left-hand side and the other on the right-side of the screen.
  • “Four Objects”: Refers to tracking part 1 (the jaws) and part 2 (the wrist) of each of the two Needle Drivers-one located on the left-hand side and the other on the right-side of the screen.
  • “Six Objects”: Refers to tracking part 1 (the jaws), part 2 (the wrist), and part 3 (the shaft) of each of the two Needle Drivers-one located on the left-hand side and the other on the right-side of the screen.
We track three parts of the tool rather than the whole tool, which captures the detailed movement of the robotic tools that could improve the accuracy of detection of key frames.
In addition, the image is divided into two halves by a vertical line at the center. At any given moment in the video, a tool is considered the left-hand tool if its centroid is located on the left half, and the right-hand tool if its centroid is located on the right half.
@@ IV-C Hyper-parameter
Our proposed methods were developed using a PyTorch framework and implemented on a Windows (R) 11 Enterprise system equipped with an AMD RyzenTM ThreadripperTM PRO 3995WX 64-cores CPU @ 2.70 GHz. We trained our model with two Nvidia RTX A6000 GPUs.
The experiments were performed based on the configurations specified in Table III. The hyperparameters such as image size, batch size, and clip length were optimized to make full use of GPU memory. Please note that the threshold values for KAFR are determined based on the desired number of extracted frames. Instead of selecting a value between 0 and 1 as in the Equation 1, we choose a number of frames by a percentage of total number of training samples, and adjust the threshold accordingly.
@@ IV-D Evaluation Metrics
[…]
@@
[…]
  3. 3.
@@ V-A Adaptive 1
In this experiment, we initially focused on tracking velocity (displacement over time) using one object—specifically, the right-hand tool, as most procedures were performed by right-hand dominant surgeons. However, this method frequently fell short of surpassing the baseline accuracy of $0.749$. This “baseline” represents the result obtained using the research settings described in prior work (See Section VI). The turning point occurred when we allocated $15\%$ of our data, resulting in an improvement that exceeded the baseline by $0.38\%$ (See Figure 9 and also Table IV).
Building upon these promising findings, we expanded our experiments to incorporate velocity tracking from two objects. The result was remarkable; with just a $10\%$ data allocation, we achieved a $2.53\%$ increase in accuracy. Our observations indicated that the method was most effective with sufficient data reduction ($20\%$, $15\%$, and $10\%$ of total frames), suggesting that critical frames were being effectively captured. Though, significant reductions in data ($5\%$ and $1\%$ of total frames) led to decreased accuracy, likely due to insufficient data for robust model training.
Subsequent studies involving additional objects yielded slight accuracy gains ($30\%$, $20\%$, $10\%$, and $5\%$), and in some cases, resulted in reduced accuracy ($50\%$, $15\%$, and $1\%$ data allocations). Though, our efforts achieved the best accuracy of $0.7684$ by utilizing four objects with a $10\%$ frame allocation.
These outcomes indicate the important role of adaptively selecting key frames in optimizing model performance. Our Adaptive 1 method, which selects frames based on recognition of significant tool movements, not only improves accuracy but also notably reduces the dataset size. In the next section, we will compare our approach with a well-known method to highlight its competitiveness.
@@ V-B Comparison
This experiment offers a comparative analysis between our proposed method, KAFR, and the widely used MSE metric. MSE, commonly employed in tasks such as detection of moving object and loss function measurement, evaluates similarity using all pixel-wise information within an image. In contrast, KAFR focuses only on specific pixels (the centroids of the tracked tools), disregarding other pixel data. This distinction highlights a key difference between the two methods: MSE provides a holistic pixel-based analysis, while KAFR concentrates exclusively on the identified tools.
Figure 10 illustrates the performance of KAFR and MSE relative to a baseline, indicating whether their accuracy has improved or decreased. KAFR uses the best results obtained from combinations of different objects, as described in the previous section. In contrast, MSE uses the same portions of the dataset but processes all pixel data. As we can see, MSE occasionally outperforms the baseline with $20\%$, $15\%$, and $5\%$ data allocations but exhibits notable limitations with $50\%$, $30\%$, $10\%$, and $1\%$ data allocations, implying vulnerability to background noise interference that impacts its performance. It should be emphasized that in surgical videos, the background which typically consists of deformable organs and connective tissues is constantly moving, making it harder to extract useful information from tool movements. On the other hand, KAFR outperforms the baseline in most cases, achieving improvements of up to approximately $2.59\%$ (X3D CNN ($10\%$)). This demonstrates its effectiveness in selectively tracking and utilizing tool data to identify key frames. In addition to comparing to the baseline, when both methods are put together, KAFR outperforms the MSE.
Figure 10: KAFR vs. MSE.
@@ V-C Adaptive 2
An alternative method to detect tool movement is to analyze the variations of the rate of change of velocity over time, referred as acceleration. These accelerations may be a sign of significant surgical maneuvers made by surgeons. Looking at the Figure 11, a trend emerges similar to that of velocity (Figure 9), where in most cases, the accuracy outperforms the baseline. However, the accuracy decreased when the number of frames is reduced to $5\%$ and $1\%$. Although the significance of accelerations in detecting crucial frames may not match that of velocity, we achieved better accuracy ($0.771$) using a $15\%$ data allocation for accelerations.
@@ V-D Two-stream Convolutional Networks
Table V shows the results achieved using the two best candidates from the previous section, namely Acceleration with Two Objects using $15\%$ of the data and Velocity with Four Objects using $10\%$ of the data. The outputs of the two channels (Optical Flow and RGB) are ensembled. As we can see, the accuracy is further improved to $0.7814$ ($4.32\%$).
@@
Accuracy
Type
X3D CNN Two Objects
Acceleration ($15\%$)
X3D CNN Four Objects
@@
Optical Flow 0.7291 0.7175
RGB 0.7438 0.7532
Ensemble 0.7814 0.7666
Figure 12: Phase Segmentation. Two distinct videos are sampled to show KAFR performance. The color codes are: ’1.1 Stay suture’ - ’Blue’, ’1.2 Inner running suture’ - ’Yellow’, ’1.3 Enterotomy’ - ’Green’, ’2.2 Inner running suture’ - ’Orange’, ’3.1 Inner Layer of Connell’ - ’Indigo’, ’4.1 Outer Layer of Connell’ - ’Violet’, ’Idle time’ - ’Gray’, respectively. Filtered Data refers to data composed of key frames.
Figure 12 illustrates the results (X3D CNN Two Objects Acceleration ($15\%$)) using two distinct videos. As shown in the figure, six phases are segmented along with idle time (in gray). The frame indices for the original data (showing gaps between frames resulting from undersampling and oversampling techniques, as more samples are extracted for shorter phases and fewer samples for longer phases) and the filtered indices after using KAFR are displayed. In addition, Table VI provides details on the total number of frames and the percentage of frames filtered for these samples. The table also shows the number of duplicated frames resulting from upsampling.
TABLE VI: Total frames and percentage of frames after filtering. The phases are: Phase 1 - ’1.1 Stay suture’, Phase 2 - ’1.2 Inner running suture’, Phase 3 - ’1.3 Enterotomy’, Phase 4 - ’2.2 Inner running suture’, Phase 5 - ’3.1 Inner Layer of Connell’, Phase 6 - ’4.1 Outer Layer of Connell’, Phase 0 - ’Idle time’, respectively. ’Dup’ is the abbreviation for duplication (number of repeated frames).
{tblr}
cells = c, row1 = Silver, row2 = Silver, row3 = Silver, cell11 = r=3, cell12 = c=4, cell22 = c=2, cell24 = c=2, vlines, hline1,4-12 = -, hline2-3 = 2-5, Phase Video ID
GJ02092021CL GJ05262020PH
Total Frames Filtered Total Frames Filtered
Phase 0 248 $11.69\%$ 248 $15.32\%$
Phase 1 250 $9.20\%$ 250 $4.00\%$
@@
Phase 3 250 $7.60\%$ 250 $12.00\%$
Phase 4 250 $24.40\%$ 250 $18.80\%$
Phase 5 250 $17.20\%$ 250 $17.20\%$
Phase 6 227 $16.30\%$ 250 $21.20\%$
Dup 138 144
The KAFR tends to skip frames at the beginning and end of a phase, e.g., the first frames in gray color and the end frames in pink color in the upper plot of Figure 12. It is also observed that in the Enterotomy phase of the upper figure, the gap for filtered data appears longer. This occurs because, in this specific instance, the surgeon performed the cutting on the right hand (the tool was not tracked), while left hand (the tool was tracked) moved more slowly or, at times, disappeared from the screen. As a result, fewer key frames were detected compared to other phases.
VI Extension
In this section, we apply our proposed methods, Adaptive 1 and Adaptive 2, to the PJ dataset (See Section IV-A2). For consistency, we maintained the same model architecture (X3D CNN) and optimized settings as detailed in previous research [3]. With this setup, the best model achieved an accuracy of 0.8801, which we use as the baseline for comparison with our findings.
Figure 13 presents a comparison of the results obtained by our proposed methods against the baseline. In this evaluation, we examined the effect of varying the training set size, using increments of $1\%$, $5\%$, $10\%$, $15\%$, $20\%$, $30\%$, and $50\%$ of the total training data. Our study also explored the impact of tracking data from multiple objects on model accuracy. Specifically, we analyzed scenarios where tracking was applied using one, two, and four objects in the same manner as with the GJ dataset.
Our best results outperformed the baseline, demonstrating the effectiveness of our adaptive methods. The highest accuracy for Adaptive 1 reached $0.8909$ (a $1.22\%$ increase) when using velocity data from one object with $30\%$ of the data. For Adaptive 2, the top accuracy was $0.8976$ (a $1.98\%$ increase), achieved with acceleration data from two objects and $20\%$ of the data. The accuracy is increased to $0.8982$ ($2.05\%$) with the ensembling method, employing Adaptive 2 with two objects approach on $20\%$ of the data.
VII State-of-the-Art Methods
TABLE VII: State-Of-The-Art Methods on Distinct Datasets.
[…]
Cholec80 TeCNO [56] 88.56
Endonet + HHMM [31] 81.70
LoViT [57] 92.40
ST-ERFNet [58] 86.07
MTRCNet-CL [58] 87.40
Cholec51 TeCNO [56] 87.34
CATARACTS SurgPLAN [59] 83.10
M2F + GCN [60] 76.86
M2cai16 LAST [61] 91.50
PJ X3D CNN [3] 88.01
@@ VII State-of-the-Art Methods
As seen on Table VII, the TeCNO represents a Multi-Stage Temporal Convolutional Network (MS-TCN) that uses hierarchical prediction refinement approach for surgical phase recognition. It employs causal, dilated convolutions, to provide a large receptive field, enabling smooth, online predictions even during ambiguous transitions in surgical phases. In contrast to standard causal MS-TCNs, the TeCNO restricts each prediction to rely only on current and past frames, enhancing real-time application by reducing dependent on future frames. Applying this method to the Cholec80 and Cholec51 datasets yielded accuracies of $88.56\%$ and $87.34\%$, respectively. Although TeCNO arguably surpasses many state-of-the-art LSTM approaches, training ResNet50 frame-by-frame without temporal context may reduce effectiveness by loosing crucial frame-to-frame relationships.
@@
The Long Video Transformer (LoViT) takes another approach by focusing the extraction of temporal spatial features and modeling phase transitions. It utilizes a multiscale temporal aggregator that integrates L-Trans modules based on self-attention to capture short-term, fine-grained details, and a G-Informer module utilizing ProbSparse self-attention to examine long-term temporal dynamics. These features are fused through a multiscale temporal head, which employs phase transition-aware supervision for accurate classification. The model achieves an accuracy of $92.40\%$ on the Cholec80 dataset. However, it does not take into consideration the dynamics of surgical tool movement, such as velocity and acceleration, which are addressed in Adaptive 1 and Adaptive 2. Additionally, the source code was unavailable at the time of writing this manuscript, preventing a fair comparison by applying the LoViT model to our datasets.

### v3

@@ III-B2 Two-stream Convolutional Networks
Back to the Figure 1, in the Phase Segmentation module, we trained two distinct X3D models separately. In the last layer of both models, instead of outputting classifications, we modified the layer to output probabilities. We then ensembled these probabilities to make classifications.
@@ IV-A Benchmark Dataset
Pancreatic cancer is projected to become the second leading cause of cancer-related deaths in the USA within the next two to three decades [mizrahi2020pancreatic]. Robotic Pancreaticoduodenectomy (also known as the Whipple procedure) surgery, which is a key procedure for treating pancreatic cancer, involves two main phases: Resection and Reconstruction. The Reconstruction phase includes three anastomoses: Hepaticojejunostomy, Pancreaticojejunostomy, and Gastrojejunostomy.
@@ IV-A1 GJ Dataset
A retrospective case review spanning from 2017 to 2021 was conducted at two prominent referral centers, UT Southwestern and University of Pittsburgh Medical Center. During this review, each robotic GJ video was segmented into six specific tasks: 1.1 Stay suture, 1.2 Inner running suture, 1.3 Enterotomy, 2.2 Inner running suture, 3.1 Inner Layer of Connell, and 4.1 Outer Layer of Connell, along with idle time. These videos were annotated by experts using a custom annotation framework (See Figure 6.
[ [..484848removed: tbhp! ] hbtp!]
[..494949removed: Render X3D model architecture. (A) The X3D network extends a 2-D ResNet (res) one axis at a time across various axes, including temporal duration $X_{t}$, frame rate $X_{T}$, spatial resolution $X_{s}$, width $X_{w}$, bottleneck width $X_{b}$, and depth $X_{d}$. (B) Details of the 2-D ResNet demonstrate residual learning (Adapted from [feichtenhofer2020x3d] and [he2016deep]). ] Refer to caption
@@
IV-B [..505050removed: Architecture ]
[..525252removed: Figure 1 presents an overview of our proposed architecture for KAFR. This workflow is divided into three distinct steps: Object Tracking, KAFR, and Phase Segmentation. In the first phase, we extract the bounding box, class ID (with a confidence score greater than $0.5$) ] Of the 42 videos included, 30 were used for training and 12 for validation, with all frames extracted (six frames/second) and annotated, resulting in a dataset of 46, [..535353removed: and frame number from the input video for surgical instruments. In the second phase, we compute the location or centroid of a tool, enabling us to track its movement. The distance and velocity are utilized as indicators to detect whether the surgeon’s maneuvers are essential. Subsequently, critical frames are identified. Afterwards, we employ X3D CNN to classify frames into six different phases. ] 903 images for training and 22,647 images for validation.
IV-A1 PJ Dataset
A retrospective review was conducted on recording videos collected from 2011 to 2022 across two quaternary referral centers: UT Southwestern Medical Center and the University of Pittsburgh Medical Center [al2024development]. Out of the 100 videos reviewed, 60 were utilized for model training, 10 for hyperparameter optimization, and 30 for performance testing. Frames were extracted at 6 frames/second and annotated. Each video was divided into tasks, with start and end times noted for each. Six tasks were annotated per video: 1.1 Anterior Mattress Sutures, 1.2 Tying Mattress Sutures, 2.1 Enterotomy, 3.1 Posterior Duct Sutures, 3.2 Anterior Duct Sutures, and 4.1 Anterior Buttress Sutures.
IV-B Distribution of Tools
Figure 7 displays the distribution of the number of frames for each class ID corresponding to the sixteen objects listed in Table II. We used YOLOV8 [redmon2016you], as described in Section III-B, to extract and track surgical instruments. Each instrument is segmented into three parts, each assigned a unique ID: the shaft, wrist, and jaw, except for the “laptool”. It’s important to note that in our dataset, instrument number two (“instrument2”) is identical to instrument number three (“instrument3”) because they are both needle drivers; therefore, they are combined under the same object category.
Figure 7: [..575757removed: Proposed Architecture ] Distribution of [..585858removed: KAFR ] Class IDs. Total of sixteen objects are tracked.
@@ IV-C Hyper-parameter
[..595959removed: The experiments in this research ] Our proposed methods were developed using a PyTorch framework and implemented on a Windows (R) 11 Enterprise system equipped with an AMD RyzenTM ThreadripperTM PRO 3995WX 64-cores CPU @ 2.70 GHz. We trained our model with two Nvidia RTX A6000 GPUs.
The experiments were performed based on the configurations specified in Tables III. The hyperparameters such as image size, batch size, and clip length were optimized to make full use of GPU memory. Please note that the threshold values for KAFR are determined based on the desired number of extracted frames. Instead of selecting a value between 0 and 1 as in the Equation 1, we choose a number of frames by a percentage of total number of training samples, and adjust the threshold accordingly.
@@ IV-D Evaluation Metrics
    Accuracy: it computes the percentage of correct predictions over all predictions made on the test data; that is, it represents the proportion of predictions that were correctly classified relative to the total number of predictions.
    Accuracy=Number of correct predictionsTotal number of $\text{Accuracy}=\frac{\text{Number of correct predictions}}{\text{Total numberof predictions}}$ (9)
@@
    F1 score: it establishes a uniform metric for evaluating the precision and recall by taking into account both false positives and false negatives.
@@ V-A Adaptive 1
In this experiment, we initially focused on tracking velocity using a single tool—specifically, the right-hand tool, as most people are comfortable with their right hand. However, this method frequently fell short of surpassing the baseline accuracy of $0.749$ (During our experiments, we identified the optimal parameter that yielded the most favorable results and this “baseline” represents a best finding or achievement derived from our research efforts). The turning point occurred when we allocated 15% of our data to the task, resulting in a substantial improvement that exceeded the baseline almost $0.22\%$ (See Figure 8 and also Table IV). Encouraged by this accomplishment, we increased our efforts to refine our approach further.
Building upon these promising findings, we expanded our experiments to incorporate velocity tracking from two tools (left and right hand). The result was remarkable; with just a 10% data allocation, we achieved a $1.9$% increase in accuracy. Our observations indicated that the method was most effective with sufficient data reduction (20%, 15%, and 10% of total frames), suggesting that critical frames were being effectively captured. Though, significant reductions in data (5% and 1% of total frames) led to decreased accuracy, likely due to insufficient data for robust model training.
Subsequent studies involving additional tools yielding slight accuracy gains (30%, 20%, 10%, and 5%), and in some cases, resulting in reduced accuracy (50%, 15%, and 1% data allocations). Though, our efforts achieved in a better accuracy of $0.7684$ by utilizing four tools with a 10% frame allocation.
These outcomes indicate the important role of adaptive learning in optimizing model performance. Our Adaptive 1 method, which selects frames based on recognition of significant tool movements, not only improves accuracy but also notably reduces the dataset size. In the next section, we will compare our approach with a well-known method to highlight its superiority.
@@ V-B Comparison
This experiment provides a comparative analysis between our proposed method, Kinematics Adaptive Frame Recognition (KAFR), and Mean Squared Error (MSE). MSE, a widely used metric in tasks such as detection of moving object and loss function measurement, evaluate changes over time using all pixel-wise information within an image. In contrast, KAFR focuses just on the tracked tools, disregarding other pixel data. This distinction highlights a key difference between the two methods: MSE considers the entirety of pixel information, while KAFR concentrates solely on the identified tools.
Figure 9 illustrates the performance of KAFR and MSE relative to a baseline, indicating whether their accuracy has improved or decreased. MSE occasionally outperforms the baseline with 20%, 15%, and 5% data allocations but exhibits notable defeats with 50%, 30%, 10%, and 1% data allocations, implying vulnerability to background noise interference that impacts its performance. It should be emphasized that in surgical videos, the background which typically consists of deformable organs and connective tissues is constantly moving, making it harder to extract useful information from tool movements. On the other hand, KAFR outperforms the baseline by approximately $1.94$%, demonstrating effectiveness in selectively tracking and utilizing tool data to identify key frames. In addition to comparing to the baseline, when both methods are put together, KAFR performs better than MSE.
Figure 9: KARF vs. MSE.
@@ V-C Adaptive 2
An alternative method to detect tool movement is to examine speed of changes in velocity, called acceleration. These accelerations may be a sign of significant surgical maneuvers made by the surgeons. Looking at the Figure 10, a trend emerges similar to that of velocity, where in most cases, the accuracy outperforms the baseline. However, the accuracy reduced when the number of frames is decreased to 5% and 1%. Although the significance of accelerations in detecting crucial frames may not match that of velocity, we achieved better accuracy ($0.771$) using a 15% data allocation for accelerations.
@@ V-D Two-stream Convolutional Networks
[..606060removed: We integrated optical flow (OF) into our approach for video action recognition because of its critical role in capturing motion dynamics and temporal changes. Optical flow estimation involves computing angle and magnitude shifts for each pixel across consecutive video frames. The Farneback [farneback2003two] algorithm is a common technique for dense optical flow computation, that utilizes polynomial expansion and iterative field estimation. By solving a quadratic polynomial equation between two consecutive frames $I(x,y,t)$ and $I(x+\Delta x,y+\Delta y,t+\Delta t)$, optical flow $(u,v)$ at each pixel is estimated to obtain displacement fields. ]
[..616161removed: In our work, we implemented the Farneback optical flow estimation using the calcOpticalFlowFarneback function from OpenCV. We set the number of pyramid levels to $3$, the pyramid scale to $0.5$, the window size to $15$, the number of iterations to $3$, the pixel neighborhood size to $5$, and a standard deviation of $1.2$. The process includes converting frames to gray-scale, using the calcOpticalFlowFarneback function to compute flow vectors, and saving the magnitude as an image. The estimated optical flow may be utilized as input to our X3D CNN model. ]
Table V shows the results achieved using the two best candidates from the previous section, namely Acceleration with 2-Tool using 10% of the data and Velocity with 4-Tool using 15% of the data. The outputs of the two channels (Optical Flow and RGB) are ensembled. As we can see, the accuracy is further improved to $0.7814$.
@@
{tblr}
row1 = Silver, row2 = Silver, column3 = c, column5 = c, cell11 = r=2, cell12 = c=2c, cell14 = c=2c, cell32 = c, cell34 = c, cell42 = c, cell44 = c, cell52 = c, cell54 = c, vlines, hline1,3-6 = -, hline2 = 2-5, Type X3D CNN Two Objects
Acceleration ($15\%$) X3D CNN Four Objects
@@
Accuracy F1 Score Accuracy F1 Score
Optical Flow 0.7291 0.6496 0.7175 0.6304
RGB 0.7438 0.6747 0.7532 0.6750
Ensemble 0.7814 0.7141 0.7666 0.6794
Figure 11: Phase Segmentation. Two distinct videos are sampled to show KAFR performance. The color codes are: ’1.1 Stay suture’ - ’Blue’, ’1.2 Inner running suture’ - ’Yellow’, ’1.3 Enterotomy’ - ’Green’, ’2.2 Inner running suture’ - ’Orange’, ’3.1 Inner Layer of Connell’ - ’Indigo’, ’4.1 Outer Layer of Connell’ - ’Violet’, ’Idle time’ - ’Gray’, respectively.
Figure 11 illustrates the results (X3D CNN 2Tool Acceleration (10%)) using two distinct videos. As shown in the figure, six phases are segmented along with idle time (in gray color). The frame indices for the original data (showing gaps between frames resulting from undersampling and oversampling techniques, as more samples are extracted for shorter phases and fewer samples for longer phases) and the filtered indices after using KAFR are displayed. The KAFR tends to skip frames at the beginning and end of a phase, e.g., the first frames in gray color and the end frames in pink color in the upper figure. It is also observed that in the Enterotomy phase of the upper figure, the gap for filtered data appears longer. This is because, in this specific instance, the surgeon performed the cutting more slowly compared to other tasks in distinct phases.
VI Extension
In this section, we apply our proposed methods, Adaptive 1 and Adaptive 2, to the PJ dataset (See Section IV-A1). For consistency, we maintained the same model architecture (X3D CNN) and optimized settings as detailed in previous research [al2024development]. With this setup, the best model achieved an accuracy of 0.8801, which we use as the baseline for comparison with our findings.
VII [..626262removed: Discussion ]
[..646464removed: Our ] Figure 12 presents a comparison of the results obtained by our proposed methods against the baseline. In this evaluation, we examined the effect of varying the training set size, using increments of $1\%$, $5\%$, $10\%$, $15\%$, $20\%$, $30\%$, and $50\%$ of the training data. Our study also explored the impact of tracking data from multiple tools on model accuracy. Specifically, we analyzed scenarios where tracking was applied using one, two, and four tools in the same manner as with the GJ dataset.
Our best results outperformed the baseline, demonstrating the effectiveness of our adaptive methods. The highest accuracy for Adaptive 1 reached 0.8909 (a $1.01\%$ increase) when using velocity data from one tool with 30% of the data. For Adaptive 2, the top accuracy was 0.8976 (a $1.75\%$ increase), achieved with acceleration data from two tools and 20% of the data. The accuracy is increased to 0.8982 with the ensembling method, employing Adaptive 2 with 2-Tool approach on 20% of the data.
VII Comparison with the state of the art
Previously, our group conducted research on the PJ dataset using the X3D CNN model. In this study, we build upon that work by introducing a new method, Kinematics Adaptive Frame Recognition, to selectively capture key frames rather than using all frames. This approach reduces the number of frames by fivefold and improves accuracy by approximately 2%.
TABLE VI: State-Of-The-Art Methods on Distinct Datasets.
@@ VII Comparison with the state of the art
Cholec80 TeCNO [czempiel2020tecno] 88.56
Endonet + HHMM [twinanda2016endonet] 81.70
LoViT [liu2025lovit] 92.40
ST-ERFNet [jin2020multi] 86.07
MTRCNet-CL [jin2020multi] 87.40
Cholec51 TeCNO [czempiel2020tecno] 87.34
CATARACTS SurgPLAN [luo2024surgplan] 83.10
M2F + GCN [holm2023dynamic] 76.86
M2cai16 LAST [tao2023last] 91.50
PJ X3D CNN [al2024development] 88.01
@@
The TeCNO represents a Multi-Stage Temporal Convolutional Network (MS-TCN) that uses hierarchical prediction refinement approach for surgical phase recognition. It employs causal, dilated convolutions, to provide a large receptive field, enabling smooth, online predictions even during ambiguous transitions in surgical phases. In contrast to standard acausal MS-TCNs, the TeCNO restricts each prediction to rely only on current and past frames, enhancing real-time application by reducing dependent on future frames. Applying this method to the Cholec80 and Cholec51 datasets yielded accuracies of 88.56% and 87.34%, respectively. Although TeCNO arguably surpasses many state-of-the-art LSTM approaches, training ResNet50 frame-by-frame without temporal context may reduce effectiveness by loosing crucial frame-to-frame relationships. Notably, its accuracy is worse than that of our KAFR approach on the PJ dataset.
@@
The Long Video Transformer (LoViT) takes another approach by focusing the extraction of temporal spatial features and modeling phase transitions. It utilizes a multiscale temporal aggregator that integrates L-Trans modules based on self-attention to capture short-term, fine-grained details, and a G-Informer module utilizing ProbSparse self-attention to examine long-term temporal dynamics. These features are fused through a multiscale temporal head, which employs phase transition-aware supervision for accurate classification. The model achieves an accuracy of 92.40% on the Cholec80 dataset, slightly outperforming our result on the PJ dataset (89.82%). However, it does not take into consideration the dynamics of surgical tool movement, such as velocity and acceleration, which are addressed in Adaptive 1 and Adaptive 2. Additionally, the source code was unavailable at the time of writing this manuscript, preventing a fair comparison by applying the LoViT model to our datasets.